OpenStudy (anonymous):

solve by elimination or substitution: 0.1x+3.2y=2 2x+64y=20

OpenStudy (anonymous):

This means that you can either use: .1x+3.2y=2 2x+64y=20 -2x-64y=-40 { *(-20) 2x+64y=20 0 = -20 => false Or: .1x+3.2y=2 2x+64y=20 <=> x=10-32y .1*(10-32y)+3.2y=2 1-3.2y+3.2y=2 1=2 => false

OpenStudy (anonymous):

so there is no solution for this question?

OpenStudy (anonymous):

Yes. There are no possible real values for x and y.

OpenStudy (anonymous):

which means that the 2 lines that are given are parallel to each other, so they will not intersect at any point.

OpenStudy (anonymous):

thank you! brilliant.